# Smaller and More Flexible Cuckoo Filters

### v2

@@ Abstract
Cuckoo filters are space-efficient approximate set membership data structures with a controllable false positive rate (FPR) and no false negatives, similar to Bloom filters. In contrast to Bloom filters, Cuckoo filters store multi-bit fingerprints of keys in a hash table using variants of Cuckoo hashing, allowing each fingerprint to be stored at a small number of possible locations. Existing Cuckoo filters use fingerprints of $(k+3)$ bits per key and an additional space overhead factor of at least $1.05$ to achieve an FPR of $2^{-k}$. For $k=10$, this amounts to 1.365kn1.365\,kn bits to store $n$ keys, which is better than 1.443kn1.443\,kn bits for Bloom filters. The +3+3 for the fingerprint size is required to balance out the multiplied FPR caused by searching for the fingerprint at $2^{3}=8$ locations. In the original Cuckoo filter, the number of hash table buckets is restricted to a power of 2, which may lead to much larger space overheads, up to 2.1(1+3/k)kn2.1\,(1+3/k)\,kn bits.
We present two improvements of Cuckoo filters. First, we remove the restriction that the number of buckets must be a power of 2 by using a different placement strategy. Second, we reduce the space overhead factor of Cuckoo filters to 1.06(1+2/k)1.06\,(1+2/k) by using overlapping windows instead of disjoint buckets to maintain the load threshold of the hash table, while reducing the number of alternative slots where any fingerprint may be found. The is parallelized implementation using a producer-consumer method which uses lock-free communication via message buffers.
@@
Our implementation of bucketed and windowed Cuckoo filters, and the workflows to reproduce all results are available at https://gitlab.com/rahmannlab/cuckoo-filters.
@@ 1 Introduction
are space efficient data structures for approximate fast set membership queries. They have many practical applications, such as for database systems [25], networks [15, 26], storage systems [22] and sequence analysis in computational biology [18, 21, 24]. A probabilistic filter represents a set $K$ of keys from a key universe $\mathcal{U}$ and supports at least two operations: (1) inserting a key, and (2) querying if a key is contained in the set. Instead of storing an exact representation of the set $K$, the keys are fuzzily represented in a way that queries for keys that were inserted into the filter are always correctly answered (no false negative results). Queries about keys that were not inserted are correctly answered most of the time, but sometimes incorrectly positively. The probability of obtaining a false positive answer is called the false positive rate (FPR). The FPR is controlled by the user and related to the space requirements of the filter. To represent an arbitrary set of keys $K$ with cardinality $|K|=n$ with an FPR of $\varepsilon=2^{-k}$, the theoretically optimal space requirement is $nk$ bits. Practical implementations of filters need $Cnk$ bits with an overhead factor $C>1$ that differs between filter types. Different filter types try to achieve small $C$ and fast insertion and query times; often with different trade-offs between space and time efficiency. In addition, some filter variations support additional operations, such as deletion of keys (dynamic filters), counting the number of times a key was inserted (counting filters), or even storing arbitrary values with the keys (probabilistic key-value stores). Other variants optimize the space requirements for a fixed set of keys, where the complete set of keys has to be known before filter construction starts (static filters).
Table 1: Overview of non-extendable filter types supporting online insertions. The overhead factor $C=C(k)>1$ specifies the required space per key; one needs CnkCnk bits to store $n$ keys with an FPR of $2^{-k}$. If a filter only supports a number of slots that is a power of $2$, $C(k)$ can vary between the given formula (best case) and twice that value (worst case; here given as worst-case $C(8)$ for a typical $k=8$ scenario). The number of cache misses for insertion/lookup is a proxy for insertion/lookup time.
@@
However, alternative filter designs often perform better than Bloom filters. Cuckoo filters [12], Morton filters [3], (Vector) Quotient filters [14, 29] and Prefix filters [11] store a $k$-bit fingerprint in a hash table instead of distributed bits in a bit array. This design allows for storing additional values together with the key fingerprints, simply by using additional bits, giving us not only a probabilistic set membership data structure, but also a probabilistic key-value store without additional effort. The FPR is controlled by the fingerprint size, where larger fingerprints lead to smaller FPRs. Different filter types in this class differ in their hash collision resolution strategies. For example, Cuckoo filters resolve collisions using Cuckoo hashing [28, 12] and Vector Quotient filters use ideas from Robin Hood hashing [29]. Depending on the collision resolution strategy, the filters achieve different load thresholds. For a hash table with $s$ slots, the load (or fill rate) $r$ is given by $r:=n/s$, where $n$ is the number of inserted keys. The load threshold (or maximum load factor) is the largest load $r$ for which the filter can be successfully constructed with high probability. To achieve low space overhead we need to design a data structure that has a high load threshold while at the same time limiting the number of possible positions a key’s fingerprint may be stored, as each alternative position would increase the FPR if not compensated for by additional bits per key.
@@
In this work, we focus on fingerprint based non-extendable filters supporting online insertions; an overview is given in Table 1. In contrast to the Bloom filter, the (best possible) overhead factor of the fingerprint based filters decreases with increasing $k$ (see also Figure 1), but many of these filters have additional technical or efficiency-related restrictions, which can increase the overhead by a factor of (almost) 2, and hinders flexible use.
@@ 2.1 Multi-way Cuckoo hashing
    $(d,l)$ windowed Cuckoo hashing works similarly to $(d,l)$ bucketed Cuckoo hashing, but uses overlapping windows instead of non-overlapping buckets [20]. Each key can be inserted into one of $dl$ slots, but each window overlaps by l1l-1 positions with its surrounding windows, hence the total number of windows is W=sl+1W=s-l+1.
@@ 2.2 Cuckoo filters
Since only store the fingerprint is stored, the two bucket addresses must be restricted so that one can be calculated from the other and the fingerprint. If we do not want to sacrifice a bit that remembers whether the fingerprint is stored in its first or second bucket, we must use a symmetric function to compute both buckets. The standard Cuckoo filter solves this problem as follows [12]. A hash function $f_{1}:\mathcal{U}\to[B]$ computes the first bucket address of a key $x$, given by $b_{1}=f_{1}(x)$. The second bucket address is the bit-wise XOR of a hash of the fingerprint $h(f_{0}(x))$ and the first bucket address $b_{1}$, where $h:[2^{q}]\to[B]$. Conversely, the first bucket address is obtained back by XORing $b_{2}$ with the same fingerprint hash. The XOR operation is only guaranteed to return a valid address if the number of buckets is a power of 2. A proposed simplification [10] uses the $q$-bit fingerprint $f_{0}(x)$ directly instead of a hash value $h(f_{0}(x))$.
@@ 3.2 Windowed layout
Thus, the main differences of our proposal to the previous Cuckoo filter [12] are that here,
@@
    we reduce the number of offset bits from 2 to 1, saving space, because the load threshold for Cuckoo hashing with two slots in windows and for four slots in buckets are comparable ($0.965$ vs. $0.980$). More precisely, filling the hash tables slightly below their theoretical load thresholds, the overhead factor is reduced from $1.05(1+3/k)$ in [12] to $1.06(1+2/k)$, e.g., from $1.365$ to $1.272$ for $k=10$ (an FPR of $1/1024$).
@@ 4.3 Bit-parallel engineering
We load the contents of a window into a 64-bit integer register that we (conceptually) partition into four adjacent slots (and some unused bits). To check whether there is an empty slot (all zeros), we use two pre-computed masks, lo and hi = lo << (q-1), indicating the least significant and most significant bits in each slot, respectively. By computing the bit pattern e := (x - lo) & (˜x) & hi, the 1-bits of e are those 1-bits of hi that indicate which slots are empty. If e is non-zero, there exists an empty slot and its index (from the right) can be computed by dividing the number of trailing zeros of e (obtained by a cttz CPU instruction) by q. In the example , the result is nonzero and has 7 trailing zeros, indicating that slot 0 in $x$ is empty. Comparing all 4 slots at the same time in this bit-parallel manner saves the overhead of a loop or separate comparisons.
@@
Similarly, to search for a fingerprint in any of the four slots, we can create a single bit mask that contains the valid bit patterns for each slot. For $(2,4)$ windowed Cuckoo filters, we use one choice bit and two window offset bits. Assume that the 5-bit fingerprint is 11100, that we are examining window choice 1, and that the offset bits increase with decreasing slot number. Then the valid slot bit patterns can be encoded in a mask y, and we can check if the bit patterns agree in one of the slots (and which one) by taking the bit-wise XOR xˆy and checking for a zero slot like above. In the following example, we find the fingerprint in slot 1.
@@ 5.3 Performance Benchmarks
We evaluate insertion throughput and query throughput separately for successful and unsuccessful queries for different $k$, with a load factor of 98% of the load threshold. Figure 5 (top) shows the throughput in million keys per second (wall time) using five inserter and query threads and different values of $k$. Insertion throughput is lower compared to query throughput. This is expected since the random walk may cause several cache misses. In contrast, queries incur at most 2 cache misses. The $(2,4)$ windowed Cuckoo filters have the highest insertion throughput, and $(2,2)$ windowed Cuckoo filters have the lowest insertion throughput, which is in concordance with the results from Section 5.1. The throughput of successful and unsuccessful queries is similarly high (Figure 5B and 5C).
@@
The speedup factor from parallelization is almost linear, both for insertions and for queries for up to 5 threads (Figure 5 bottom, see Appendix 4.4 for details). For more threads, the distributing main thread starts to become the bottleneck. The speedup for insertions is slightly lower than for queries, probably due to higher memory bandwidth utilization.
@@ 5.4 Comparison with Prefix Filters and Vector Quotient Filters
We evaluated the insertion and lookup throughput (50% contained and 50% not contained keys) on random 64-bit integers (see Table 3). Prefix filters have the highest insertion and lookup throughput. Cuckoo filters have the lowest insertions throughput due to the many cache misses compared to $\approx 1$ cache miss for Prefix filters and 2 for Vector Quotient filters. For lookups, Vector Quotient filters and Cuckoo filters have similar throughput due to the same amount of cache misses. For $k=8$ Prefix filters have the highest throughput, but in the optimized cases (i.e., Cuckoo filters with 16 bits per slot), bucketed Cuckoo filters have a higher lookup throughput compared to Prefix and VQF filters.
@@
The space overhead depends on the FPR and a valid comparison is thus only possible for the same FPR. For a given FPR, the $(2,2)$ windowed Cuckoo filter has the smallest space overhead compared to Prefix, VQF and bucketed Cuckoo filters (see Table 3).
@@ 6 Conclusion and Discussion
We introduced improvements of Cuckoo filters, enabling their more flexible use and reducing their overhead factor from $1.05(1+3/k)$ to $1.06(1+2/k)$, switching from a $(2,4)$ bucketed layout to a $(2,2)$ windowed layout. This is a practical improvement for relevant FPRs. For an FPR of $2^{-8}$, $(2,2)$ windowed Cuckoo filters need $\approx 10.6$ GB to index 10 billion keys, compared to $\approx 11.6$ GB for $(2,4)$ bucketed Cuckoo filters. Another interesting aspect of this work is that just-in-time compiled Python code can adjust to run-time specified parameters (such as $k$), resulting in a flexible implementation of Cuckoo filters that is competitive in speed with a C++ implementation.

### v3

@@ Abstract
Cuckoo filters are space-efficient approximate set membership data structures with a controllable false positive rate (FPR) and no false negatives, similar to Bloom filters. In contrast to Bloom filters, Cuckoo filters store multi-bit fingerprints of keys in a hash table using variants of Cuckoo hashing, allowing each fingerprint to be stored at a small number of possible locations. Existing Cuckoo filters use fingerprints of $(k+3)$ bits per key and an additional space overhead factor of at least $1.05$ to achieve an FPR of $2^{-k}$. For $k=10$, this amounts to 1.365kn1.365\,kn bits to store $n$ keys, which is better than 1.443kn1.443\,kn bits for Bloom filters. The +3+3 for the fingerprint size is required to balance out the multiplied FPR caused by searching for the fingerprint at $2^{3}=8$ locations. In the original Cuckoo filter, the number of hash table buckets is restricted to a power of two, which may lead to much larger space overheads, up to 2.1(1+3/k)kn2.1\,(1+3/k)\,kn bits (e.g. 2.73kn2.73\,kn bits for $k=10$).
We present two improvements of Cuckoo filters. First, we remove the restriction that the number of buckets must be a power of two by using a different placement strategy. Second, we reduce the space overhead factor of Cuckoo filters to 1.06(1+2/k)1.06\,(1+2/k) (e.g. $1.272$ instead of $1.365$ for $k=10$) by using overlapping windows instead of disjoint buckets. Thereby, we maintain a comparable load threshold of the hash table, while reducing the number of alternative slots where any fingerprint may be found and decrease the fingerprint size from $k+3$ to $k+2$.
The implementation is parallelized using a producer-consumer framework which uses lock-free communication via message buffers.
@@
Our implementation of bucketed and windowed Cuckoo filters, and the workflows to reproduce all results are available at https://gitlab.com/rahmannlab/cuckoo-filters. Artifacts can also be found at zenodo, DOI https://doi.org/10.5281/zenodo.17078618.
@@ 1 Introduction
are space efficient data structures for approximate fast set membership queries. They have many practical applications, such as for database systems [26], networks [16, 27], storage systems [23] and sequence analysis in computational biology [19, 22, 25]. A probabilistic filter represents a set $K$ of keys from a key universe $\mathcal{U}$ and supports at least two operations: (1) inserting a key, and (2) querying if a key is contained in the set. Instead of storing an exact representation of the set $K$, the keys are fuzzily represented in a way that queries for keys that were inserted into the filter are always correctly answered (no false negative results). Queries about keys that were not inserted are correctly answered most of the time, but sometimes incorrectly positively. The probability of obtaining a false positive answer is called the false positive rate (FPR). The FPR is controlled by the user and related to the space requirements of the filter. To represent an arbitrary set of keys $K$ with cardinality $|K|=n$ with an FPR of $\varepsilon=2^{-k}$, the theoretically optimal space requirement is $nk$ bits. Practical implementations of filters need $Cnk$ bits with an overhead factor $C>1$ that differs between filter types. Different filter types try to achieve small $C$ and fast insertion and query times; often with different trade-offs between space and time efficiency. In addition, some filter variations support additional operations, such as deletion of keys (dynamic filters) [12, 15, 31], counting the number of times a key was inserted (counting filters) [13, 30], or even storing arbitrary values with the keys (probabilistic key-value stores) [35]. Other variants optimize the space requirements for a fixed set of keys, where the complete set of keys has to be known before filter construction starts (static filters) [17, 18, 8].
Table 1: Overview of non-extendable filter types supporting online insertions. The overhead factor $C=C(k)>1$ specifies the required space per key; one needs CnkCnk bits to store $n$ keys with an FPR of $2^{-k}$. If a filter only supports a number of slots that is a power of $2$, $C(k)$ can vary between the given formula (best case, see Figure 1) and twice that value (worst case; here given as worst-case $C(8)$ for a typical $k=8$ scenario). The number of cache misses for insertion/lookup is a proxy for insertion/lookup time.
@@
However, alternative filter designs often perform better than Bloom filters. Cuckoo filters [12], Morton filters [3], (Vector) Quotient filters [15, 31] and Prefix filters [11] store a $k$-bit fingerprint in a hash table instead of distributed bits in a bit array. This design allows for storing additional values together with the key fingerprints, simply by using additional bits, giving us not only a probabilistic set membership data structure, but also a probabilistic key-value store without additional effort. The FPR is controlled by the fingerprint size, where larger fingerprints lead to smaller FPRs. Different filter types in this class differ in their hash collision resolution strategies. For example, Cuckoo filters resolve collisions using Cuckoo hashing [29, 12] and Vector Quotient filters use ideas from Robin Hood hashing [31].
Depending on the collision resolution strategy, the filters achieve different load thresholds. For a hash table with $s$ slots, the load (or fill rate) $r$ is given by $r:=n/s$, where $n$ is the number of inserted keys. The load threshold (or maximum load factor) is the largest load $r$ for which the filter can be successfully constructed with high probability. To achieve low space overhead we need to design a data structure that has a high load threshold while at the same time limiting the number of possible positions a key’s fingerprint may be stored, as each alternative position would increase the FPR if not compensated for by additional bits per key.
@@
In this work, we focus on fingerprint-based non-extendable filters supporting online insertions; an overview is given in Table 1. In contrast to the Bloom filter, the (best possible) overhead factor of the fingerprint-based filters decreases with increasing $k$ (see also Figure 1), but many of these filters have additional technical or efficiency-related restrictions, which can increase the overhead by a factor of (almost) 2, and hinders flexible use.
@@ 2.1 Multi-way Cuckoo hashing
    $(d,l)$ windowed Cuckoo hashing works similarly to $(d,l)$ bucketed Cuckoo hashing, but uses overlapping windows instead of non-overlapping buckets [21]. Each window with $l$ slots overlaps by l1l-1 positions with its surrounding windows, hence the total number of windows is W=sl+1W=s-l+1. Everything else works analogously to the bucketed version.
@@ 2.2 Cuckoo filters
Since only the fingerprint is stored, the two bucket addresses must be restricted so that one can be calculated from the other and the fingerprint. If we do not want to sacrifice a bit that remembers whether the fingerprint is stored in its first or second bucket, we must use a symmetric function to compute both buckets. The standard Cuckoo filter solves this problem as follows [12]. A hash function $f_{1}:\mathcal{U}\to[B]$ computes the first bucket address of a key $x$, given by $b_{1}=f_{1}(x)$. The second bucket address $b_{2}$ is given by the bit-wise XOR of a hash of the fingerprint $h(f_{0}(x))$ and the first bucket address $b_{1}$, where $h:[2^{q}]\to[B]$. Conversely, the first bucket address is obtained back by XORing $b_{2}$ with the same fingerprint hash. The XOR operation is only guaranteed to return a valid address if the number of buckets is a power of 2. A proposed simplification [10] uses the $q$-bit fingerprint $f_{0}(x)$ directly instead of a hash value $h(f_{0}(x))$.
@@ 3.2 Windowed layout
In summary, we need a fingerprint size of $k+2$ in a $(2,2)$ windowed Cuckoo filter for a FPR of $2^{-k}$ and in practice use a maximum fill rate of $\approx 0.94$ (see Figure 4). The space overhead of a $(2,2)$ windowed Cuckoo filter is therefore $1/0.94(1+2/k)\approx 1.06(1+2/k)$. The main differences of our proposal to the previous Cuckoo filter [12] are that in our work,
@@
    we reduce the number of offset bits from 2 to 1, saving space, because the load threshold for Cuckoo hashing with two slots in windows and for four slots in buckets are comparable ($0.965$ vs. $0.980$). Filling the hash tables slightly below their theoretical load thresholds, the overhead factor is reduced from $1.05(1+3/k)$ in [12] to $1.06(1+2/k)$, e.g., from $1.365$ to $1.272$ for $k=10$ (an FPR of $1/1024$).
@@ 4.3 Bit-parallel engineering
We load the contents of a window into a 64-bit integer register that we (conceptually) partition into four adjacent slots (and some unused bits). To check whether there is an empty slot (all zeros), we use two pre-computed masks, lo and hi = lo << (q-1), indicating the least significant and most significant bits in each slot, respectively. By computing the bit pattern e := (x - lo) & (˜x) & hi, the 1-bits of e are those 1-bits of hi that indicate which slots are empty. If e is non-zero, there exists an empty slot and its index (from the right) can be computed by dividing the number of trailing zeros (obtained by a cttz LLVM intrinsic) of e by q. In the example , the result is nonzero and has 7 trailing zeros, indicating that slot 0 in $x$ is empty. Comparing all 4 slots at the same time in this bit-parallel manner saves the overhead of a loop or separate comparisons.
@@
Similarly, to search for a fingerprint in any of the four slots, we can create a single bit mask that contains the valid bit patterns for each slot. For $(2,4)$ windowed Cuckoo filters, we use one choice bit and two window offset bits. Assume that the 5-bit fingerprint is 11100, that we are examining window choice 1, and that the offset bits increase with decreasing slot number. Then the valid slot bit patterns can be encoded in a mask y, and we can check if the bit patterns agree in one of the slots (and which one) by taking the bit-wise XOR xˆy and checking for a zero slot like above. The two possible masks for y that encode the valid bit pattern for the first and second window choice are precomputed based on $d$ and $l$ and used as compile-time constants in the lookup function. In the following example, we find the fingerprint in slot 1.
@@ 4.4 Parallelization
Assigning the keys randomly to $F$ independent subfilters causes a randomly varying load factor per subfilter. However, this effect is negligible for large enough $n$ because the variation in the number of keys per subfilter (grows with $\sqrt{n}$) is much smaller than the extra space needed to guarantee that all keys can be inserted into the filter with high probability (grows linearly with $n$). For example, For $F=10$ subfilters and $n=10^{8}$ keys, the number of keys in each subfilter follows a Binomial distribution with success rate $p=1/10$. Hence, the expected number of keys in each subfilter is $np=10^{7}$ and the standard deviation is $\sqrt{np\,(1-p)}=3000$, but the space overhead already accounts for $0.06\cdot np=600\,000$ additional slots per subfilter. Hence, for typically large $n$, the parallelization does not incur an additional space overhead compared to a single-threaded version.
@@ 5.3 Performance Benchmarks
We evaluate insertion throughput and query throughput separately for successful and unsuccessful queries for different $k$, with a load factor of 98% of the load threshold. Figure 5 (top) shows the throughput in million keys per second (wall time) using five inserter and query threads and different values of $k$. Insertion throughput is lower compared to query throughput. This is expected since the random walk may cause several cache misses. In contrast, queries incur at most two cache misses. The $(2,4)$ windowed Cuckoo filters have the highest insertion throughput, and $(2,2)$ windowed Cuckoo filters have the lowest insertion throughput, which is in concordance with the results from Section 5.1. The throughput of successful and unsuccessful queries is similarly high (Figure 5B and 5C).
@@
The speedup factor from parallelization is almost linear, both for insertions and for queries for up to 5 threads (Figure 5 bottom, see Section 4.4 for details). For more threads, the distributing main thread starts to become the bottleneck. The speedup for insertions is slightly lower than for queries, probably due to higher memory bandwidth utilization.
@@ 5.4 Comparison with Prefix Filters and Vector Quotient Filters
We evaluated the insertion and lookup throughput (50% contained keys and 50% not contained keys) on random 64-bit integers (see Table 3). Prefix filters have the highest insertion and lookup throughput. Cuckoo filters have the lowest insertions throughput due to the many cache misses compared to $\approx 1$ cache miss for Prefix filters and two for Vector Quotient filters. For lookups, Vector Quotient filters and Cuckoo filters have similar throughput due to the same amount of cache misses. For $k=8$ Prefix filters have the highest throughput, but in the optimized cases (i.e., Cuckoo filters with 16 bits per slot), bucketed Cuckoo filters have a higher lookup throughput compared to Prefix and VQF filters. The lower throughput of windowed vs. bucketed Cuckoo filters is mainly due to the fact that the windows are not cache-aligned and a certain fraction of windows spans two cache lines.
@@
The space overhead depends on the FPR and a valid comparison is thus only possible for the same FPR. For a given FPR, the $(2,2)$ windowed Cuckoo filter has the smallest space overhead compared to Prefix, VQF and bucketed Cuckoo filters (see factors $C$ in Table 3).
@@ 6 Conclusion and Discussion
We introduced improvements of Cuckoo filters, enabling their more flexible use and reducing their overhead factor from $1.05(1+3/k)$ to $1.06(1+2/k)$, by switching from a $(2,4)$ bucketed layout to a $(2,2)$ windowed layout. This is an improvement for relevant FPRs (e.g. around $k=10$, up to $k\leq 102$). For an FPR of $2^{-8}$, $(2,2)$ windowed Cuckoo filters need $\approx 10.6$ GB to index 10 billion keys, compared to $\approx 11.6$ GB for $(2,4)$ bucketed Cuckoo filters. Another interesting aspect of this work is that just-in-time compiled Python code can adjust to run-time specified parameters (such as $k$), resulting in a flexible implementation of Cuckoo filters that is competitive in speed with a C++ implementation.